### v2

# A search for transit timing variations in the transiting hot Jupiter systems HIP 65, NGTS-6, NGTS-10 and WASP-173

Abstract
Hot Jupiters are Jupiter-mass planets with orbital periods of less than ten days. Their short orbital separations make tidal dissipation within the stellar host especially efficient, potentially leading to a measurable evolution of the orbit. One possible manifestation of this is orbital decay, which presents itself observationally through variations in the orbital period and thus times of transit. Here we select four promising exoplanetary systems for detecting this effect: HIP 65, NGTS-6, NGTS-10 and WASP-173. We present 33 new transit light curves taken with the 1.54 m Danish Telescope, and analyse these alongside photometric data from the Transiting Exoplanet Survey Satellite and transit timing data from the literature. We construct two ephemeris models for each target: a linear ephemeris and a shrinking orbital period due to tidal decay. The linear ephemeris is preferred for three of the four models – the highest significance for the quadratic ephemeris is over 3 for WASP-173. We compare these results to theoretical predictions for tidal dissipation of gravity waves in radiation zones, and find that wave breaking is predicted only in WASP-173, making rapid decay plausible in this system but unclear in the other three. The sensitivity of transit timings to orbital decay depends on the square of the time interval covered by available observations, so our results establish a useful baseline against which future measurements can be compared. NGTS-6 and NGTS-10 are important objects for future study as they are in the first field to be observed by the upcoming PLATO mission.
keywords:
methods: data analysis – techniques: photometric – planetary systems.1 Introduction
Hot Jupiters, classified based on their large masses and short orbital periods, are a relatively rare type of exoplanet within the underlying population (2010Sci...330..653H; 2012ApJS..201...15H; mayor2011harpssearchsouthernextrasolar). Despite their scarcity, they make up a significant fraction of the known exoplanets (Gaudi_2002; 2012ApJ...753..160W). This is a consequence of their relatively large sizes and small semimajor axes making them the easiest transiting planets to detect. This proximity to the host star raises tides on both bodies, causing tidal bulges (Murray_Dermott_2000).
For a planet in an asynchronous orbit – one in which the planet’s spin period is not equal to the orbital period – the motion of the tidal bulges is opposed by viscous friction. This dissipation of energy transfers angular momentum from the planetary orbit to the stellar spin. In the typical case for a hot Jupiter, where its orbital period is shorter than the host star’s spin period, tidal effects spin up the star and shrink the orbit (1980A&A....92..167H; 1981A&A....99..126H). The eventual outcome of sustained orbital shrinkage is the planet slowly spiralling into the star in a process called tidal decay (1973ApJ...180..307C). It is worth noting that there are various contributions to tidal dissipation (1996ApJ...470.1187R), but arguably the largest is that of the wave-like dynamical tide (2020MNRAS.498.2270B).
Tidal decay in exoplanets is not easily detectable. There is currently no confirmed way to measure the spin or internal density distribution of exoplanets, which would otherwise give information on the tidal mechanisms and angular momentum transfer within the system. However, some claims have been made for the observation of oblateness brought on by tides (e.g. 2024ApJ...976L..14L; 2025ApJ...981L...7P; 2025ApJ...983..157Z). The timescales on which tides act are also unknown and it has been argued that the circularization of transiting exoplanets may not be entirely due to tides (2009ApJ...692L...9L).
Presently, signs of tidal decay have only been detected through transit-timing variations (TTVs; 2018haex.bookE...7A), which show variations in the times of mid-transit and a decrease in the orbital period of the planet over time. The associated change in period is measurable from transit light curves and allows for the determination of the modified tidal quality factor, , providing an insight into the rate of energy dissipation within the star due to tides. is defined as follows:
where is the Love number (1911spge.book.....L) and is the tidal quality factor – linked to the dissipation of tidal energy through the equation (1966Icar....5..375G):
where is the maximum energy stored in a tide, is the dissipation rate and the integral is over one tidal cycle. While a negative period derivative can be a response to the tidal evolution of the system, this effect can be caused by other phenomena such as the light-time effect, in which a wide-orbiting third body causes the star-planet pair to orbit the system’s barycentre (1952ApJ...116..211I), and apsidal motion, arising only in orbits with non-zero eccentricity. Both phenomena are periodic so can be distinguished from tidal decay over a large enough timescale.
WASP-12 b, a 1.5 MJup hot Jupiter orbiting an F-type main-sequence star (Hebb_2009), was until recently the sole confirmed case of tidal decay in exoplanets (Maciejewski+16aa; 2020ApJ...888L...5Y; 2021AJ....161...72T). Since the confirmation of the system’s shrinking orbital period, the decay rate has been refined multiple times (2022MNRAS.512.3113B; 2022AJ....163..281T; 2022AJ....163..175W; 2023TJAA....4...10A; 2023JAVSO..51..243N; 2024A&A...685A..63A; 2024A&A...686A..84L; 2024ApJS..270...14W). A recent study by 2025A&A...694A.233B has detected orbital decay within WASP-43, making it the second confirmed hot Jupiter system. Despite the contemporary exclusivity of these systems, there are many candidates for tidal decay, with few having established TTVs.
One such candidate is WASP-4 b (2008ApJ...675L.113W), for which TTVs were first detected by 2019AJ....157..217B. These authors suggested the apparent period change could be caused by tidal decay, apsidal precession or gravitational perturbation from another body. 2019MNRAS.490.4230S refined the rate of period shrinkage and ruled out stellar activity and the Applegate mechanism (1992ApJ...385..621A). 2020ApJ...893L..29B and 2023Univ....9..506H investigated the possibility of a line-of-sight acceleration, attributed to a distant perturbing body. With their new transit timing data, 2025MNRAS.541..714B established that the discrepancies would be best explained by tidal orbital decay. However, the TTVs of the WASP-4 system are now believed to be associated with the light-time effect (2025arXiv251005229W).
1.1 Target selection
The magnitude of detectable orbital decay within hot Jupiter planetary systems can be quantified using the transit time shift equations presented by 2014MNRAS.440.1470B and 2018AcA....68..371M:
where is the shift in transit midpoints over ten years, and are the respective stellar and planetary masses, is the stellar radius, is the orbital semimajor axis and is the orbital period. Here, is set to a canonical value of (2014ARA&A..52..171O). Whilst requiring some preliminary spectroscopic parameters, this equation allows systems with the most favourable predicted time shifts to be selected for analysis.
We used this equation to select targets with the largest . We then restricted the sample based on sky position to include only those objects with good observability for the available telescope time. A nominal lower limit of 10 s was specified for , but objects close to this limit were ultimately not observed due to the limited availability of telescope time. The first result of this project was an analysis of HATS-18 (2022MNRAS.515.3212S); see also a study of KELT-16 by 2022MNRAS.509.1447M.
1.1.1 HIP 65 Ab
Our first target is HIP 65, a wide binary star system hosting a K-dwarf (HIP 65 A) and an M-dwarf (HIP 65 B). The hot-Jupiter planet HIP 65 Ab orbits the K-dwarf with an ultra-short period of 0.981 d (Nielsen_2020; 2021AJ....162..176P). HIP 65 A and HIP 65 B are separated by 3.95 arcsec on the sky (Nielsen_2020), corresponding to a projected separation of 269 au. These authors highlighted the planet’s large impact parameter and grazing eclipse, making radius estimations unreliable. They stated also that HIP 65 Ab is one of very few massive planets that reside within twice their Roche limit, lying on the border of the Neptune desert (Szabó_2011). However, its exact relation to the Roche limit is uncertain due to the grazing eclipses. Nielsen_2020 also report HIP 65 A’s rotation period as d from rotational modulation in light curves, and suggest tidal spin-up due to the discrepancy between the gyrochronological ( Gyr) and isochronal ( Gyr) ages of the star. However, the isochronal ages of K-dwarfs are unreliable so this is not in itself evidence of tidal evolution (Maxted++15aa2; but see also 2021ApJ...919..138T).
Photometric data on HIP 65 Ab from the Transiting Exoplanet Survey Satellite (TESS; 2015JATIS...1a4003R) have previously been examined for eclipse depth variations (2022sas..conf....7G; 2024AJ....167....1W), with the purpose of tracking its atmospheric activity. No eclipse depth variations were found after fixing the impact parameter; as such, the grazing eclipse was likely responsible for the initial variations.
Using equation 3, we find HIP 65 Ab to be a promising candidate for detecting tidal decay, having an estimated of seconds in 10 years (s in 10 yr). In spite of the promise of this system, there have been two previous orbital decay searches of HIP 65 Ab with limited success. 2024PSJ.....5..163A found a period increase of using 106 transit times spanning five years. 2024A&A...692A..35M found the decay rate to be days per orbital cycle. Both studies were limited by the data available at the time, with the majority of their transit midpoints originating from TESS. Here, we use new precise transit photometry alongside TESS to search for variations in transit timing.
1.1.2 NGTS-6 b
Our second target is an ultrashort-period hot Jupiter orbiting an old K-dwarf. NGTS-6 b (2019MNRAS.489.4125V) is the least massive planet in our sample, 1.3 MJup, and has an orbital period of 0.882 d. The estimated orbital decay associated with this planet is in 10 yr. 2020MNRAS.498.2270B suggested that the stellar host likely experiences no wave breaking, so rapid decay is not expected. However, photometric observations may provide signs of tidal decay in the event that its internal gravity waves are fully damped. This exoplanet also orbits near Roche contact (2022ApJ...939...91A) similar to the case of HIP 65 Ab.
1.1.3 NGTS-10 b
NGTS-10 b is a 2.16 MJup planet with an ultrashort period of 0.767 days and had the shortest known period of any hot Jupiter at the time of discovery (2020MNRAS.493..126M). Its stellar host is an old K-type star of mass 0.70 M⊙, which has a rotation period of 17.3 d. There have been various theoretical studies of the system, some investigating its atmosphere and climate (e.g. 2023A+A...671A.122H; 2024ApJ...975...61K), and some investigating the effects of tides (e.g. 2020MNRAS.498.2270B; 2021MNRAS.506.2247A; 2024ApJ...973..128T). NGTS-10 b is an interesting object for several research areas, patticularly orbital decay. We estimate a large of in 10 yr, but as with NGTS-6, no wave breaking or rapid decay is projected (2020MNRAS.498.2270B).
1.1.4 WASP-173 Ab
WASP-173 Ab (KELT-22 A) is a hot Jupiter orbiting a solar-type star, and is notable for being discovered by two independent consortia simultaneously (2019MNRAS.482.1379H; 2019ApJS..240...13L). Similarly to the HIP 65 system, its host star has a distant binary companion, this time at 1400 au. WASP-173 Ab is an inflated planet, with a mass of 3.7 MJup, a radius of 1.2 RJup, a period of 1.39 d, and an irradiation level significantly more than the threshold for inflation (2011ApJS..197...12D; 2019ApJS..240...13L). Its stellar host is a G-type star with a rotation period of around (2024A&A...690A.379K).
This planet has seen little individual attention since its discovery, mostly being included in tide-related projects with substantial sample sizes (e.g. 2021ApJ...919..138T). The linear ephemeris has been refined numerous times in previous large-scale studies (e.g. 2022ApJS..259...62I; 2023ApJS..264...37S; 2023ApJS..265....4K) and analysed for timing variations (2024PSJ.....5..163A; 2024A&A...692A..35M), with many suggesting it is a good candidate for tidal decay (e.g. 2020AJ....159..150P; 2024ApJ...960...50W). With the time shift equation, we estimate a period decay of over the next 10 years and contrast it with the increase in orbital period found by 2024PSJ.....5..163A of , and that of days per cycle by 2024A&A...692A..35M. We aim to refine these prior studies with the addition of our new transit light curves, whilst also providing data for future work in this area.
2 Observations
2.1 Danish 1.54-m Telescope
| Target | Date of first | Start time | End time | (s) | (s) | Filter | Airmass | Moon | Aperture | |
| observation | (UT) | (UT) | illumination | (pixels) | ||||||
| HIP 65 | 2021/09/20 | 04:32 | 07:17 | 230 | 30 | 11 | 1.11 1.26 | 0.992 | 11 20 40 | |
| HIP 65 | 2021/09/22 | 04:06 | 06:31 | 210 | 30 | 11 | 1.11 1.19 | 0.983 | 25 35 70 | |
| HIP 65 | 2022/09/15 | 05:17 | 06:46 | 75 | 60 | 12 | 1.11 1.17 | 0.747 | 20 32 70 | |
| HIP 65 | 2022/09/16 | 04:58 | 07:24 | 114 | 60 | 12 | 1.11 1.25 | 0.658 | 17 25 45 | |
| HIP 65 | 2022/09/21 | 02:51 | 05:14 | 165 | 40 | 12 | 1.19 1.11 | 0.212 | 22 32 70 | |
| HIP 65 | 2023/07/15 | 08:26 | 10:28 | 101 | 60 | 13 | 1.12 1.11 1.14 | 0.054 | 25 40 80 | |
| HIP 65 | 2023/07/17 | 07:21 | 09:48 | 122 | 60 | 13 | 1.17 1.11 1.12 | 0.003 | 25 35 80 | |
| HIP 65 | 2023/09/13 | 04:32 | 06:55 | 196 | 30 | 10 | 1.12 1.11 1.17 | 0.032 | 20 30 60 | |
| HIP 65 | 2024/07/18 | 04:31 | 06:50 | 261 | 20 | 9 | 1.67 1.21 | 0.878 | 20 30 60 | |
| HIP 65 | 2024/07/19 | 04:08 | 06:09 | 229 | 20 | 11 | 1.78 1.28 | 0.938 | 20 30 60 | |
| HIP 65 | 2024/09/04 | 06:21 | 08:15 | 420 | 5 | 10 | 1.11 1.25 | 0.014 | 12 24 40 | |
| NGTS-6 | 2019/08/23 | 07:11 | 10:26 | 235 | 40 | 10 | 1.90 1.04 | 0.524 | 8 18 35 | |
| NGTS-6 | 2019/09/07 | 07:53 | 10:01 | 80 | 80 | 11 | 1.24 1.03 | 0.630 | 19 28 50 | |
| NGTS-6 | 2019/09/15 | 07:24 | 09:58 | 125 | 60 | 14 | 1.23 1.00 1.24 | 0.986 | 8 20 50 | |
| NGTS-6 | 2021/09/25 | 05:11 | 07:15 | 156 | 30 | 15 | 1.65 1.14 | 0.832 | 7 20 40 | |
| NGTS-6 | 2021/10/02 | 06:12 | 08:29 | 187 | 30 | 13 | 1.24 1.01 | 0.202 | 8 19 40 | |
| NGTS-6 | 2022/10/11 | 06:05 | 08:00 | 97 | 35–60 | 14 | 1.16 1.01 | 0.977 | 7 18 35 | |
| NGTS-6 | 2023/10/05 | 05:54 | 07:53 | 148 | 30 | 14 | 1.26 1.02 | 0.667 | 13 22 50 | |
| NGTS-6 | 2024/09/05 | 07:48 | 09:59 | 175 | 30 | 13 | 1.27 1.02 | 0.047 | 9 19 40 | |
| NGTS-6 | 2024/09/13 | 06:27 | 08:20 | 178 | 30–20 | 13 | 1.51 1.10 | 0.711 | 8 20 40 | |
| NGTS-10 | 2022/09/12 | 07:17 | 09:52 | 103 | 100–55 | 12 | 1.71 1.07 | 0.951 | 8 12 50 | |
| NGTS-10 | 2022/09/22 | 06:39 | 09:08 | 79 | 100 | 14 | 1.70 1.08 | 0.128 | 9 25 40 | |
| NGTS-10 | 2023/10/09 | 04:36 | 08:27 | 125 | 100 | 13 | 2.46 1.05 | 0.256 | 15 22 40 | |
| NGTS-10 | 2023/10/12 | 06:32 | 08:28 | 63 | 100 | 11 | 1.28 1.03 | 0.056 | 9 12 17 | |
| NGTS-10 | 2024/09/18 | 07:10 | 09:32 | 85 | 60 | 13 | 1.57 1.06 | 0.999 | 6 12 50 | |
| NGTS-10 | 2024/09/28 | 06:21 | 09:38 | 13 | 60 | 13 | 1.64 1.01 | 0.172 | 8 25 40 | |
| WASP-173 | 2021/09/29 | 03:03 | 07:24 | 390 | 20 | 14 | 1.03 1.00 1.44 | 0.490 | 21 32 70 | |
| WASP-173 | 2022/07/06 | 05:19 | 08:42 | 244 | 30 | 17 | 1.59 1.01 | 0.419 | 7 38 52 | |
| WASP-173 | 2022/09/05 | 05:14 | 09:12 | 543 | 10–20 | 13 | 1.00 1.51 | 0.676 | 8 30 50 | |
| WASP-173 | 2023/08/01 | 06:02 | 10:16 | 835 | 5–4 | 13 | 1.08 1.00 1.20 | 0.995 | 8 25 50 | |
| WASP-173 | 2023/08/26 | 04:49 | 09:15 | 522 | 10 | 17 | 1.05 1.00 1.32 | 0.701 | 8 35 50 | |
| WASP-173 | 2024/07/03 | 04:52 | 09:22 | 317 | 30 | 13 | 1.91 1.00 | 0.083 | 8 30 50 | |
| WASP-173 | 2024/08/22 | 02:52 | 07:05 | 732 | 10–5 | 13 | 1.37 1.00 1.02 | 0.486 | 9 30 50 |
All new ground-based observations in this work were obtained using the 1.54 m Danish Telescope at ESO La Silla, Chile, as a side-project of the MiNDSTEp microlensing observations (Dominik+10an). The Danish Faint Object Spectrograph and Camera (DFOSC) was used in imaging mode, in which form it is equipped with a 20482048 pixel CCD camera with a 13.7′ field of view sampled at 0.39′′ px-1. The CCD was windowed to decrease the readout time in all cases. The filters used were Johnson , Bessell and Bessell . Some of the observations were obtained with the telescope moderately defocussed to improve the precision of the observations, following the method set out by Me+09mn. An observing log is given in Table 1. All observations in the log were included in the timing analysis.
HIP 65 is a bright star () and the planet produces a short (47 min) and shallow (0.6%) eclipse. We experimented with using different filters and focus levels, in some cases pre-planned and in others as a response to the sky conditions during an observing sequence. The majority of our observations were, in the end, obtained through the filter and with moderate defocus, and achieved a photometric precision in the region of 1 mmag per point.
NGTS-6 is rather fainter () and has a companion at an angular distance of 5.4′′. The majority of our observations were taken with no or small defocus, and through the filter to lower the count rate. When the seeing was good we were able to extract light curves of NGTS-6 without contamination by the fainter nearby star.
NGTS-10 is another relatively faint star () but without a nearby companion. In all our observing sequences we used an filter to maximise throughput, and moderate defocussing with exposure times of 60–100 s. The photometric precisions obtained are in the region of 1.5 mmag per point.
WASP-173 is a bright star () with a fainter companion at 6.0′′. When the observing conditions were good we used short exposure times and operated the telescope in focus to extract light curves of WASP-173 without contamination. In times of poorer seeing we defocussed the telescope and increased the exposure time, with the intention of recording light curves of the combined light of WASP-173 and its companion. This means some of our light curves accurately reflect the properties of WASP-173 itself, whereas some suffer from third light which causes a smaller transit depth.
The data were reduced using the defot pipeline (Me+09mn2; Me+14mn), which in turn uses the idl111https://www.ittvis.com/idl/ implementation of the aper routine from daophot (Stetson87pasp) contained in the NASA astrolib library222http://idlastro.gsfc.nasa.gov/ to perform aperture photometry. We constructed master bias and flat-field images but did not apply them to the data because their main effect was to add to the scatter in the light curves without modifying the shape of the transit. A differential-magnitude light curve was generated for each transit observation by constructing an optimal composite comparison star to calculate differential magnitudes against. The composite comparison star was made by iteratively adjusting the weights of individual stars and the coefficients of a low-order polynomial to minimise the scatter in the data outside transit.
The timestamps for the midpoint of each image were taken from the headers of the fits files and converted to the BJDTDB timescale using routines from Eastman++10pasp. Manual time checks were performed for many transits, in all cases confirming the reliability of the timestamps in the fits headers.
2.2 Transiting Exoplanet Survey Satellite
TESS data were extracted from the Mikulski Archive for Space Telescopes (MAST) archive333https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html via the Lightkurve package (2018ascl.soft12013L). We exclusively used 120 s cadence observations from SPOC, with a standard bitmask. Pre-search Data Conditioning Simple Aperture Photometry (PDCSAP) fluxes (2016SPIE.9913E..3EJ) were used where possible. The first and final 1% of times in each light curve were discarded to remove any artifacts. We applied the Savitzky-Golay filter (1964AnaCh..36.1627S) to remove low-frequency trends before the light curves were normalized. Fluxes were then converted to magnitude units. Each TESS sector was split into two, where the most central transit midpoint was taken from each half-TESS sector for the light curve and timing analyses.
HIP 65 A has been observed in six TESS sectors (1, 2, 28, 29, 68 and 69) spanning a five-year time interval, with our ground-based transit light curves adding an additional year. WASP-173 A has been observed in half as many sectors (2, 29 and 69), NGTS-6 was only observed in two sectors (5 and 32) and NGTS-10 has coverage from just one sector (87).
3 Analysis
3.1 Light curve analysis
We used version 43 of the jktebop code (2013A&A...557A.119S) to fit all light curves. Free parameters for our ground-based observations include the sum of fractional radii (, where ), ratio of the radii (), inclination (), light scale factor and time of mid-transit (). The orbital period () was also made a free parameter during light curve fitting of TESS data. We applied the power-2 limb darkening law (1997A&A...327..199H), with coefficients interpolated from tabulations by 2022A&A...664A.128C. These coefficients were kept fixed. We assume fixed circular orbits for all targets based on the adopted values established in all respective discovery papers, despite some small but non-zero eccentricity results from other sources (e.g. 2023ApJS..265....4K). All other initial and fixed transit parameters were taken from the respective discovery papers. Uncertainties on fitted parameters were calculated using both Monte Carlo and residual-shift methods in jktebop (2008MNRAS.386.1644S). The residual-shift method assesses the importance of correlated noise, accounting for it in the error budget. The overall uncertainties on the mid-transit times were picked as the larger of the two multiplied by the corresponding reduced value. Fig. 1 shows all new transit light curves from the Danish Telescope.
3.2 Literature mid-times
We combined our transit times for HIP 65 Ab with an additional three from the literature. These come from the ExoClock Project (2022ExA....53..547K; 2022ApJS..258...40K; 2023ApJS..265....4K), an operation aiming to track the ephemerides of over 1000 transiting exoplanets. The three transit times were taken in succession with the Yves Jongen Telescope in Deep Sky Chile. Three transit times were also pulled from the Exoplanet Transit Database (ETD), a catalogue of transit light curves taken by amateur astronomers (2010NewA...15..297P).
Eight transit times of NGTS-6 b are available from the ExoClock Project: four from Yves Jongen at Deep Sky Chile, three from Jean-Pascal Vignes at Deep Sky Chile and one from Anaël Wünsche at El Sauce Observatory. These times span the month of December in the year 2021. We refitted one light curve using the NGTS data published in the planet’s discovery paper (2019MNRAS.489.4125V) and three additional times were taken from the ETD. We note that six of the transit midpoints from ExoClock have an epoch in common with an ExoClock transit time. Since all common times are taken with different telescopes, we do not reject any of the data. The same applies to a transit time from the ETD which was obtained at the same epoch as one we observed with the Danish Telescope.
Available timing data for the NGTS-10 system is scarce. The only source in the literature is the discovery paper (2020MNRAS.493..126M), from which we obtain four transit midpoints by refitting the transit light curves. The resulting times were converted from HJDUTC to BJDTDB following the method described in Section 2.1. No data are currently available from ExoClock or the ETD.
In the case of WASP-173 Ab, one transit light curve was taken from the first discovery paper by 2019MNRAS.482.1379H and refitted. Timing data were given in HJDUTC and so were converted to BJDTDB. Three mid-transit times were presented in the second discovery paper (2019ApJS..240...13L). The ExoClock Project and ETD again both have timing data. We obtained one transit time from each.
3.3 Transit timing analysis
| HIP 65 Ab | ||
| Quantity | Linear model | Quadratic model |
| (BJD) | 2459540.547764(42) | 2459540.547697(64) |
| (d) | 0.980972186(48) | 0.980972242(63) |
| (d) | - | |
| 27 | 26 | |
| 49.9 | 48.0 | |
| BIC | 60.0 | 61.5 |
| AIC | 55.9 | 56.0 |
| log | - | |
| NGTS-6 b | ||
| Quantity | Linear model | Quadratic model |
| (BJD) | 2459550.679388(94) | 2459550.67962(14) |
| (d) | 0.88205815(11) | 0.88205809(11) |
| (d) | - | |
| 23 | 22 | |
| 66.0 | 60.5 | |
| BIC | 75.7 | 73.4 |
| AIC | 72.0 | 68.5 |
| log | - | |
| NGTS-10 b | ||
| Quantity | Linear model | Quadratic model |
| (BJD) | 2460035.788622(83) | 2460035.78876(10) |
| (d) | 0.766893317(50) | 0.76689297(17) |
| (d) | - | |
| 10 | 9 | |
| 17.6 | 13.0 | |
| BIC | 25.0 | 23.0 |
| AIC | 23.6 | 21.0 |
| log | - | |
| WASP-173 Ab | ||
| Quantity | Linear model | Quadratic model |
| (BJD) | 2459486.704505(71) | 2459486.704763(98) |
| (d) | 1.38665315(11) | 1.38665272(15) |
| (d) | - | |
| 17 | 16 | |
| 58.0 | 43.5 | |
| BIC | 66.8 | 55.3 |
| AIC | 64.0 | 51.5 |
| log | - | |
We constructed linear and quadratic ephemerides to fit the timing data. The linear fit represents a constant-period orbit in a system with no orbital decay. It is described by
where is the constant orbital period, is the transit epoch and is the value at . The quadratic fit represents an orbital decay model, but is also sensitive to the other aforementioned TTV mechanisms. This model is
and invokes a quadratic coefficient , a quantity linked to the rate of period growth by . A negative value of indicates period shrinkage and possible tidal decay. Model parameters for both fits were estimated using an MCMC method with the emcee package (2013ascl.soft03002F), where the quadratic coefficient was allowed to be positive or negative. For each model, we let 500 walkers move 20,000 steps, with an arbitrary 1000 step burn-in. We do not fit for periodic trends in this study due to the sparse data coverage.
The difference in strength between the two models was quantified by employing both the Bayesian Information Criterion (BIC; 10.1214/aos/1176344136) and Akaike Information Criterion (AIC; 1100705). These excel in testing the strength of the fits to the data, whilst also penalising overly-complex higher-order polynomials. The AIC and BIC were computed with
and
for each model, where is the polynomial degree, is the number of data points and was calculated from each model using
similarly to equation 2 in 2020AJ....159..150P. A list of all statistical quantities from the fits is displayed in Table 2, and linear residuals are plotted against the best fit in Fig. 2. We compared the two models by subtracting the quadratic BIC and AIC from their linear counterparts. Thus, a significant positive BIC and AIC should imply the presence of TTVs. The range of values of BIC and AIC that indicate a significant preference for one model over another is hard to define, so we required and to accept the quadratic model in preference to the linear one (e.g. Kass01061995).
3.4 Tidal quality factor
We determined a lower bound on the modified tidal quality factor as an indication of the efficiency of tidal dissipation within the star, and its influence over the planet’s orbit. Based on results from 2014MNRAS.440.1470B, 2018AcA....68..371M and others, and the formulation by 1966Icar....5..375G, we use the following equation for the modified tidal quality factor:
where is the rate of change of the orbital period. All other parameters have their former meanings. The lower limit on for each stellar host was obtained by implementing the 3 lower uncertainty of into the above equation, given the relationship . Limits on for the four objects are shown in Table 2. Note that this tidal quality factor represents only the stellar tidal dissipation and is independent of any dissipation within the planet.
3.5 Theoretical predictions
Our lower bound constraints on for each of the four systems can be compared to theoretical predictions. The most efficient tidal mechanism in each case is predicted to be dissipation of internal gravity waves (part of the dynamical tide response) in their radiative cores, which are launched and propagate inwards from the radiative/convective boundary. If these waves are fully damped (by whatever mechanism) we can readily predict the resulting tidal dissipation and hence (e.g. using Eq. 41 in 2020MNRAS.498.2270B). This regime is likely to be relevant if the planetary mass is large enough to cause the waves to break in the stellar core (e.g. 2010MNRAS.404.1849B), though there are other possibilities, including gradual radiative damping of the waves leading to spin-up of the central portions of the star and the subsequent efficient wave absorption (e.g. 2023MNRAS.521.1353G), other strong wave-wave interactions (e.g. 2024ApJ...960...50W), or sufficiently strong magnetic fields (though this is expected to be more relevant for F-stars; e.g. 2024ApJ...966L..14D).
We have computed MESA stellar models and the resulting tidal gravity wave for each of these systems (NGTS-6b and 10b were previously presented in 2020MNRAS.498.2270B). For HIP 65 A we used a mass and rotation period of d (Nielsen_2020), and for WASP-173 A we used a mass (2019MNRAS.482.1379H) and rotation period of d (2024A&A...690A.379K) (both with initial metallicity ). We find (at age 4.1 Gyr; Nielsen_2020) to be applicable for HIP 65 Ab’s orbital decay, and (at ages 3-7 Gyr, with smaller values for older stars) for WASP-173 Ab. Both NGTS-6b and NGTS-10b are predicted to have relevant for their orbital decay at ages of approximately 10 Gyr (2020MNRAS.498.2270B). These predictions assume the gravity waves to be fully damped, and much larger values () would be expected if the waves are only weakly damped by radiative diffusion (unless the system happens to excite a g-mode in resonance, which has a low probability). Wave breaking is predicted for the expected ages of WASP-173 A (i.e. at 3-7 Gyr, with 2019MNRAS.482.1379H finding Gyr and 2019ApJS..240...13L finding Gyr), so this regime is plausibly justified in this system, though it is less clear in the others, none of which clearly satisfy the criterion for wave breaking (in §3.2 of 2020MNRAS.498.2270B).
4 Results and Discussion
A transit timing analysis was carried out on four hot Jupiter systems: HIP 65 A, NGTS-6, NGTS-10, WASP-173 A. Their large mass ratios, short periods and large fractional radii made these planets excellent candidates for the detection of tidal decay. We collected a total of 33 transit times of our own from ground-based observations using the 1.54m Danish Telescope, 24 from TESS (two per sector), 21 from the literature and seven from the ETD. The time baselines covered by all data used in this study are , , and for HIP 65 Ab, NGTS-6b, NGTS-10b and WASP-173 Ab respectively. The results from our analyses are tabulated in Table 2 and displayed visually in Fig. 2.
4.1 HIP 65 Ab
From the orbital decay model, we find a positive quadratic coefficient, corresponding to a period increase of . For comparison, the published period derivatives for HIP 65 Ab are from 2024PSJ.....5..163A and a rate of period change per cycle of by 2024A&A...692A..35M. Our result is consistent with these but also with a constant period at under 2. The statistical analysis shows no substantial favouring of either model, with and , acting as sufficient reason not to explore models with higher-order polynomials. The detection of any orbital decay in this system would require observations over a longer time span than currently available.
4.2 NGTS-6 b
We find the difference between the strength of the linear and quadratic models to be and , which does not significantly favour the quadratic model. Alongside an inspiral rate of , tidal decay is not significantly favoured over the constant-period alternative. Although the best-fitting value is negative, this result is consistent with zero at 2.3 and therefore does not directly imply a shrinking orbit. Further observations will be paramount in distinguishing orbital decay from typical orbital motion. Opportunely, NGTS-6 b is in a sample of known transiting exoplanets that will be observed by the PLATO mission (Rauer+25exa) in its first long-pointing field (LOPS2; 2025A&A...694A.313N). As a result, the occasion may present itself in less than a decade with the release of PLATO data. This planet makes an interesting tidal decay candidate for future study.
4.3 NGTS-10 b
A successful transit timing analysis of NGTS-10 is hindered by the lack of available transit times. Although the times used in this study span over eight years, only 12 transit times have been obtained. The value presented here is negative with a significance of just above 2 based on the available data. There is a gap of several years between the earlier times and the observations presented in the current work (see Fig. 2), meaning that a linear ephemeris is precisely anchored but that the quadratic term is not well constrained.
Without precise points in the epoch space between the original discovery observations and the first Danish Telescope observation, this decay rate could vary significantly in either direction. The status of orbital decay in this system will be helped by new timing data, particularly as this planet is also in the PLATO LOPS2 field (2025A&A...694A.313N) and has planned observations with the JWST (2006SSRv..123..485G; 2024AJ....167..195C). Although theoretical studies provide conflicting evidence for the rate of tidal dissipation (e.g. 2020MNRAS.498.2270B; 2024ApJ...973..128T), longer photometric coverage will be beneficial in constraining even should it be consistent with zero. As of now, NGTS-10 b still remains a good candidate.
4.4 WASP-173 Ab
Our timing analysis of WASP-173 Ab produced a lower and than those detected in other works, with 2024PSJ.....5..163A finding and 2024A&A...692A..35M finding . The differences are likely attributed to the variety of data sources used in each work. We find and , indicating the quadratic model fits the residuals better than that of the linear. Although is negative at 3.8, this result could be optimistic, caused by underestimated uncertainties on some of the transit times. The orbital decay scenario is still not conclusive. Nevertheless, 2024A&A...692A..35M predict wave breaking provided the stellar host is older than , giving the theoretical constraint which agrees with our own theoretical estimation. As with the other three systems, more timing data spanning a longer time interval will be needed to identify orbital decay in WASP-173 A.
5 Conclusions
By taking the 3 lower limits of for each system, we computed lower limits for . These are not remarkable, approximately lying in the range , but are compatible with our theoretical predictions if tidally-driven orbital decay due to gravity waves is actually occurring. Given that the current constraints are not very far from predictions, each of these 4 systems would be promising for future follow-up studies since they have very good potential to test tidal theory. On the other hand, the possible outward migration of HIP 65 Ab, if confirmed in future studies, would be more difficult to explain by tidal processes alone in such a slowly rotating star. A significant positive , if found, would be most likely explained by another phenomenon, such as the light time effect (e.g. 2025ApJ...986..117Y) or an acceleration of the system away from the observer. As of now, our TTV models show no conclusive evidence for orbital decay in any of these systems. WASP-173 Ab has the strongest suggestion of orbital decay and is a promising candidate for future studies.
The data we presented and analysed in the current work will be valuable for future studies of these objects. Long-term monitoring of all four planetary systems will be useful in progressively constraining and ultimately leading to a significant detection of tidal effects.
Acknowledgements
JS acknowledges support from STFC under grant number ST/Y002563/1. FA, MA, and UGJ acknowledge funding from the Novo Nordisk Foundation Interdisciplinary Synergy Programme grant no. NNF19OC0057374. VB and PR are supported by PRIN 2022 CUP D53D23002590006. RFJ acknowledges the support provided by the GEMINI/ANID project under grant number 32240028, by ANID’s Millennium Science Initiative through grant ICN12_009, awarded to the Millennium Institute of Astrophysics (MAS), and by ANID’s Basal project FB210003. EK is supported by the National Research Foundation of Korea 2021M3F7A1082056. LM acknowledges the financial contribution from PRIN MUR 2022 project 2022J4H55R. This paper includes data collected by the TESS mission and obtained from the MAST data archive at the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by the NASA’s Science Mission Directorate. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. This work has made use of data from the European Space Agency (ESA) mission Gaia444https://www.cosmos.esa.int/gaia, processed by the Gaia Data Processing and Analysis Consortium DPAC555https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has received funding from the Europlanet 2024 Research Infrastructure (RI) programme. The Europlanet 2024 RI provides free access to the world’s largest collection of planetary simulation and analysis facilities, data services and tools, a ground-based observational network and programme of community support activities. Europlanet 2024 RI has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 871149.
The following resources were used in the course of this work: the NASA Astrophysics Data System; the SIMBAD database operated at CDS, Strasbourg, France; and the ariv scientific paper preprint service operated by Cornell University.
This work is based on data collected by MiNDSTEp with the Danish 1.54 m telescope at the ESO La Silla Observatory in Chile. The operation, servicing, and maintenance of the DK-1.54m telescope is supported by a Villum Young Investigator grant (project number 25501) and a Villum Experiment grant (VIL69896) from VILLUM FONDEN.
Data Availability
The light curves obtained with the Danish Telescopes will be made available at the Centre de Données astronomiques de Strasbourg (CDS) at http://cdsweb.u-strasbg.fr/. The TESS data used in this article are available in the MAST archive (https://mast.stsci.edu/portal/Mashup/Clients /Mast/Portal.html).
Appendix A Transit light curves from TESS
Appendix B Timing data used in this work
TableTTs.csv
TableTTs_ngts6_.csv
TableTTs_ngts10_.csv
TableTTs_wasp173_.csv